# Invisible Hydrodynamic Tweezer Based on Near-Zero Index Materials

### v1

@@ I Introduction
Manipulating individual cells or tissue clusters is of significant importance in life science research, particularly in stem cell research [1], regenerative medicine [2], organ-on-a-chip technology [3], and tissue engineering [4]. Since Ashkin first captured and manipulated bacteria and living cells in 1987 [5, 6], various contactless tweezers have been widely applied in these fields, contributing to breakthroughs in microbiology, molecular biology, biophysics, and bioanalytical chemistry. Contactless tweezers utilize various physical forces to capture and manipulate particles. The most well-known optical tweezers [7, 8, 9, 10] use the radiation pressure and gradient forces of a laser beam to create an optical potential well, capturing and manipulating microscopic particles. Magnetic tweezers [11] employ magnetic fields to exert forces on magnetic particles, precisely controlling their position and movement. Acoustic tweezers [12, 13] use the acoustic radiation force generated by sound waves propagating through a medium, achieving precise particle manipulation by adjusting the phase and intensity of the sound waves. Other types include optoelectronic tweezers [14], which utilize dielectrophoresis effects on photoconductive materials to manipulate particles, and plasmonic tweezers [15], which use locally enhanced electromagnetic fields generated by surface plasmon resonance to capture particles.
@@
Despite the significant advancements in the development of various types of tweezers, several limitations persist. First, regarding the manipulated objects, some tweezers may cause damage during use. For instance, the high-power lasers in optical tweezers can inflict thermal damage on cells, while the enhanced electromagnetic fields in plasmonic tweezers can lead to localized heating, potentially harming biological tissues. Furthermore, many tweezers are limited to manipulating specific types of particles; for example, magnetic tweezers are restricted to magnetic particles, and optoelectronic tweezers depend on the polarization properties of the objects. Second, concerning the background environment, existing contactless tweezers cannot eliminate interference outside their functional area, which may lead to unintended disruptions. When multiple tweezers are used simultaneously, they can interfere with each other, reducing their precision. Moreover, these tweezers struggle to stably capture targets in flowing fluids. Third, the operational principles of current contactless tweezers require active control, necessitating a continuous and stable excitation source to maintain particle capture. This requirement renders the devices and platforms for these tweezers complex and energy-intensive. Although hydrodynamic tweezers [16, 17], which utilize fluid forces to manipulate particles, have been proposed to simplify equipment, they also demand high energy input and significantly impact areas outside their functional region. Recently, the advent of hydrodynamic metamaterials [18, 19, 20, 21, 22, 23, 24, 25] offers a promising solution to several existing challenges. These metamaterials have introduced a new paradigm in flow field control, enabling precise manipulation of fluids through the meticulous design of artificial structures. This approach not only enhances the flexibility and accuracy of fluid manipulation but also facilitates the development of unprecedented functional devices. For example, a hydrodynamic cloak [26, 27, 28, 29, 30, 31, 32, 33] can reroute fluid around a device and restore it to its original path, making the flow field appear unchanged regardless of the cloak’s presence, thus achieving hydrodynamic cloaking. Furthermore, Boyko et al. introduced the concept of hydrodynamic shielding [34, 35], which creates a uniform pressure distribution in the functional area, eliminating the pressure differential forces exerted by the fluid on the object. These two functionalities inspire the design of a new class of hydrodynamic tweezers. However, to date, research has yet to simultaneously achieve both hydrodynamic shielding and cloaking, and the principles of metamaterials have not yet been applied to the design of contactless tweezers, which aim to create a damage-free, non-interfering, and passive universal tool for controlling particles in flowing fluids.
@@ II.1 The governing equations in hydrodynamics
The Poiseuille model serves as a fundamental theory for understanding fluid dynamics in confined spaces, such as pipes and narrow channels. Its applications are extensive across both theoretical fluid mechanics and practical industrial processes. This model is instrumental in modeling the flow of blood through vessels, designing and evaluating fluid movements within reactors, simulating the seepage of groundwater through the porous media of soil and rock, and enhancing fluid transport efficiency in microchannels. Specifically, for Poiseuille flow occurring between two parallel plates, the velocity distribution is governed by the equation[36]:
@@ II.2 Hydrodynamic near-zero index metamaterial
The concept of zero index metamaterials first emerged in the field of photonics [37, 38], distinguished by their remarkable properties of infinite phase velocity and wavelength. By engineering artificial microstructures, it is possible to create near-zero index metamaterials (NZIM), which allow waves to exhibit a uniform phase distribution across the material. This characteristic is particularly difficult to replicate in natural materials. Such unique behavior is demonstrated in Figure 2a, highlighting the advanced capabilities of NZIMs.
@@
The concept of NZIM extends beyond wave systems to include diffusion systems as well [39]. For example, in the context of heat transfer, an NZIM can be conceptualized as a material possessing infinite thermal conductivity, which enables heat to traverse through it without any reduction in temperature. This property can be analogously linked to the steady-state Fourier heat conduction law, expressed as $\kappa\nabla^{2}T=0$. When compared with Equation (3), both exhibit remarkably similar forms, where $T$ in the heat conduction equation corresponds to the $P$ in the fluid dynamics equation. Thus, we characterize a fluid’s near-zero index metamaterial (NZIM) as a material through which, when fluid flows, the pressure variation is minimal, denoted as $\nabla p\rightarrow 0$. Additionally, from Equation (3), we observe that the term $h^{3}/12\mu$ is analogous to $\kappa$ in heat conduction systems. Therefore, while theoretically achieving a zero-index condition necessitates the parameter $h^{3}/12\mu$ approaching infinity, this scenario is impractical in real-world applications. Firstly, when $h$ becomes exceedingly large, the assumptions underlying the Poiseuille model no longer hold, rendering Equation (3) inapplicable. Secondly, significantly altering the fluid’s viscosity coefficient $\mu$ to near zero is exceedingly challenging. Consequently, alternative strategies must be explored to realize effects similar to those of NZIMs.
@@ V Discussion and conclusion
The tweezers introduced in this study are not limited to direct applications in life sciences. They can be coupled with other physical fields to provide even more functionalities. For example, integrating this technology with heat transfer mechanisms could serve as a method for controlling convective heat transfer [40, 41, 42, 43, 44, 45]. This cross-disciplinary potential significantly broadens the scope of applications, enhancing the utility and impact of invisible hydrodynamic tweezers. In addition, the concept of hydrodynamic zero-index materials proposed in this study, due to their mechanical protection properties, holds promise for a variety of applications (see Supplementary Note 15), such as shielding bridge piers from river flow impacts. Furthermore, the method of adjusting the channel height not only facilitates the realization of the hydrodynamic zero-index phenomenon but also offers a means of achieving arbitrary flow velocities. For instance, increasing the channel height can reduce the flow velocity in a specific region, while decreasing the height can increase it. This approach, when combined with hydrodynamic metamaterials, has the potential to create devices with additional functionalities, such as hydrodynamic concentrators. Finally, there is room for further improvement in this device. Currently, we achieve various functionalities by adjusting the height of the flow channels, resulting in abrupt changes in height across different regions, which can cause flow losses. A more efficient approach would be to use continuously varying heights rather than abrupt transitions. This could potentially be achieved by employing conformal mapping techniques [46, 47].

### v2

# Metafluidic Tweezers: Precise Liquid Manipulation in Flowing Fluids

@@ I Introduction
Manipulating individual cells or tissue clusters is of significant importance in life science research, particularly in stem cell research stem cell research , regenerative medicine regenerative medicine , organ-on-a-chip technology organ-on-a-chip technolog , and tissue engineering tissue engineering . Since Ashkin first captured and manipulated bacteria and living cells in 1987 Ashkin-1 ; Ashkin-2 , various contactless tweezers have been widely applied in these fields, contributing to breakthroughs in microbiology, molecular biology, biophysics, and bioanalytical chemistry. Contactless tweezers utilize various physical forces to capture and manipulate particles. The most well-known optical tweezers optical tweezers-1 ; optical tweezers-2 ; optical tweezers-3 ; optical tweezers-4 use the radiation pressure and gradient forces of a laser beam to create an optical potential well, capturing and manipulating microscopic particles. Magnetic tweezers Magnetic tweezers employ magnetic fields to exert forces on magnetic particles, precisely controlling their position and movement. Acoustic tweezers Acoustic tweezers-1 ; Acoustic tweezers-2 use the acoustic radiation force generated by sound waves propagating through a medium, achieving precise particle manipulation by adjusting the phase and intensity of the sound waves. Other types include optoelectronic tweezers optoelectronic tweezers , which utilize dielectrophoresis effects on photoconductive materials to manipulate particles, and plasmonic tweezers plasmonic tweezers , which use locally enhanced electromagnetic fields generated by surface plasmon resonance to capture particles.
@@
Despite the significant advancements in the development of various types of tweezers, several limitations persist. First, regarding the manipulated objects, some tweezers may cause damage during use. For instance, the high-power lasers in optical tweezers can inflict thermal damage on cells, while the enhanced electromagnetic fields in plasmonic tweezers can lead to localized heating, potentially harming biological tissues. Furthermore, many tweezers are limited to manipulating specific types of particles; for example, magnetic tweezers are restricted to magnetic particles, and optoelectronic tweezers depend on the polarization properties of the objects. Second, concerning the background environment, existing contactless tweezers cannot eliminate interference outside their functional area, which may lead to unintended disruptions. When multiple tweezers are used simultaneously, they can interfere with each other, reducing their precision. Moreover, these tweezers struggle to stably capture targets in flowing fluids. Third, the operational principles of current contactless tweezers require active control, necessitating a continuous and stable excitation source to maintain particle capture. This requirement renders the devices and platforms for these tweezers complex and energy-intensive. Although hydrodynamic tweezers hydrodynamic tweezers-1 ; hydrodynamic tweezers-2 , which utilize fluid forces to manipulate particles, have been proposed to simplify equipment, they also demand high energy input and significantly impact areas outside their functional region. Recently, the advent of hydrodynamic metamaterials hydrodynamic metamaterials-1 ; hydrodynamic metamaterials-2 ; hydrodynamic metamaterials-3 ; hydrodynamic metamaterials-4 ; hydrodynamic metamaterials-5 ; hydrodynamic metamaterials-6 ; hydrodynamic metamaterials-7 ; hydrodynamic metamaterials-8 offers a promising solution to several existing challenges. These metamaterials have introduced a new paradigm in flow field control, enabling precise manipulation of fluids through the meticulous design of artificial structures. This approach not only enhances the flexibility and accuracy of fluid manipulation but also facilitates the development of unprecedented functional devices. For example, a hydrodynamic cloak hydrodynamic cloak-1 ; hydrodynamic cloak-2 ; hydrodynamic cloak-3 ; hydrodynamic cloak-4 ; hydrodynamic cloak-5 ; hydrodynamic cloak-6 ; hydrodynamic cloak-7 ; hydrodynamic cloak-8 can reroute fluid around a device and restore it to its original path, making the flow field appear unchanged regardless of the cloak’s presence, thus achieving hydrodynamic cloaking. Furthermore, Boyko et al. introduced the concept of hydrodynamic shielding hydrodynamic shield-1 ; hydrodynamic shield-2 , which creates a uniform pressure distribution in the functional area, eliminating the pressure differential forces exerted by the fluid on the object. These two functionalities inspire the design of a new class of hydrodynamic tweezers. However, to date, research has yet to simultaneously achieve both hydrodynamic shielding and cloaking, and the principles of metamaterials have not yet been applied to the design of contactless tweezers, which aim to create a damage-free, non-interfering, and passive universal tool for controlling particles in flowing fluids.
@@ II.1 The governing equations in hydrodynamics
The Poiseuille model serves as a fundamental theory for understanding fluid dynamics in confined spaces, such as pipes and narrow channels. Its applications are extensive across both theoretical fluid mechanics and practical industrial processes. This model is instrumental in modeling the flow of blood through vessels, designing and evaluating fluid movements within reactors, simulating the seepage of groundwater through the porous media of soil and rock, and enhancing fluid transport efficiency in microchannels. Specifically, for Poiseuille flow occurring between two parallel plates, the velocity distribution is governed by the equationPoiseuille :
@@ II.2 Hydrodynamic near-zero index metamaterial
The concept of zero index metamaterials first emerged in the field of photonics zero photonics-1 ; zero photonics-2 , distinguished by their remarkable properties of infinite phase velocity and wavelength. By engineering artificial microstructures, it is possible to create near-zero index metamaterials (NZIM), which allow waves to exhibit a uniform phase distribution across the material. This characteristic is particularly difficult to replicate in natural materials. Such unique behavior is demonstrated in Figure 2a, highlighting the advanced capabilities of NZIMs.
@@
The concept of NZIM extends beyond wave systems to include diffusion systems as well zero thermal . For example, in the context of heat transfer, an NZIM can be conceptualized as a material possessing infinite thermal conductivity, which enables heat to traverse through it without any reduction in temperature. This property can be analogously linked to the steady-state Fourier heat conduction law, expressed as $\kappa\nabla^{2}T=0$. When compared with Equation (3), both exhibit remarkably similar forms, where $T$ in the heat conduction equation corresponds to the $P$ in the fluid dynamics equation. Thus, we characterize a fluid’s near-zero index metamaterial (NZIM) as a material through which, when fluid flows, the pressure variation is minimal, denoted as $\nabla p\rightarrow 0$. Additionally, from Equation (3), we observe that the term $h^{3}/12\mu$ is analogous to $\kappa$ in heat conduction systems. Therefore, while theoretically achieving a zero-index condition necessitates the parameter $h^{3}/12\mu$ approaching infinity, this scenario is impractical in real-world applications. Firstly, when $h$ becomes exceedingly large, the assumptions underlying the Poiseuille model no longer hold, rendering Equation (3) inapplicable. Secondly, significantly altering the fluid’s viscosity coefficient $\mu$ to near zero is exceedingly challenging. Consequently, alternative strategies must be explored to realize effects similar to those of NZIMs.
@@ V Discussion and conclusion
The tweezers introduced in this study are not limited to direct applications in life sciences. They can be coupled with other physical fields to provide even more functionalities. For example, integrating this technology with heat transfer mechanisms could serve as a method for controlling convective heat transfer multiphysics-1 ; multiphysics-2 ; multiphysics-3 ; multiphysics-4 ; multiphysics-5 ; multiphysics-6 . This cross-disciplinary potential significantly broadens the scope of applications, enhancing the utility and impact of invisible hydrodynamic tweezers. In addition, the concept of hydrodynamic zero-index materials proposed in this study, due to their mechanical protection properties, holds promise for a variety of applications (see Supplementary Note 15), such as shielding bridge piers from river flow impacts. Furthermore, the method of adjusting the channel height not only facilitates the realization of the hydrodynamic zero-index phenomenon but also offers a means of achieving arbitrary flow velocities. For instance, increasing the channel height can reduce the flow velocity in a specific region, while decreasing the height can increase it. This approach, when combined with hydrodynamic metamaterials, has the potential to create devices with additional functionalities, such as hydrodynamic concentrators. Finally, there is room for further improvement in this device. Currently, we achieve various functionalities by adjusting the height of the flow channels, resulting in abrupt changes in height across different regions, which can cause flow losses. A more efficient approach would be to use continuously varying heights rather than abrupt transitions. This could potentially be achieved by employing conformal mapping techniques conformal mapping-1 ; conformal mapping-2 .